# A Comprehensive Survey of Scientific Large Language Models and Their Applications in Scientific Discovery

### v2

@@ 6.3 Language + Vision
Biomedical text-image pairs typically come from two sources: (1) medical reports, such as chest X-rays (e.g., MIMIC-CXR Johnson et al. (2019)) and pathology reports Huang et al. (2023b); as well as (2) figure-caption pairs extracted from biomedical papers (e.g., ROCO Pelka et al. (2018) and MedICaT Subramanian et al. (2020)).
Most biomedical vision-language models exploit the CLIP architecture Radford et al. (2021), where a text encoder and an image encoder are jointly trained to map the paired text and image closer via contrastive learning (Type 3.d). The choice of the text encoder evolves from BERT Zhang et al. (2022) and GPT-2 Huang et al. (2023b) to LLaMA Wu et al. (2023) and LLaMA-2 Liu et al. (2023b), while the image encoder evolves from ResNet Huang et al. (2021) to ViT Zhang et al. (2023c) and Swin Transformer Thawkar et al. (2024). MLM, masked image modeling, and text-text/image-image contrastive learning (i.e., by creating augmented views within the language/vision modality) are sometimes adopted as auxiliary pre-training tasks. Besides CLIP, other general-domain vision-language architectures, such as LLaVA Li et al. (2023a), PaLM-E Tu et al. (2024), and Gemini Saab et al. (2024), have been explored. For instance, LLaVA-Med (Type 2.d) encodes images onto several visual tokens and prepends them to text tokens as the LLM input. Evaluation tasks of these models encompass image classification, segmentation, object detection, vision QA, text-to-image/image-to-text retrieval, and report generation, the benchmarks of which include CheXpert Irvin et al. (2019), PadChest Bustos et al. (2020), SLAKE Liu et al. (2021a), etc.
@@ 6.5 Applications in Scientific Discovery
Similarly to chemistry, LLMs can automate experiments in biological and medical research. For example, CRISPR-GPT Huang et al. (2024) augments an LLM agent with domain knowledge to enhance the design process of CRISPR-based gene-editing experiments; TrialMind Wang et al. (2024f) utilizes LLMs to extract results and synthesize clinical evidence from the literature for medical discovery. Moreover, LLMs can encode biological sequences to capture structural properties, guide protein design, and evaluate the evolutionary fitness of viral variants. For instance, ESM-1b Rives et al. (2021) and ESM-2 Lin et al. (2023b) enable accurate structure prediction of proteins without expensive and time-consuming experiments; Ferruz and Höcker (2022) fine-tune LLMs on protein families, which can generate highly divergent but still potentially functional novel sequences; Hie et al. (2021) develop LLMs that can predict viral escape mutations.
@@ 8 Challenges and Future Directions
This survey primarily covers LLMs in mathematics and natural sciences. We are aware that LLMs can also significantly impact social sciences by achieving remarkable performance in representative tasks Ziems et al. (2024) and serving as agents for social simulation experiments Horton (2023), but we leave the survey of these efforts as future work due to space limitations. In addition, this paper focuses on LLMs pre-trained on scientific data or augmented with domain-specific knowledge to benefit scientific discovery. There are studies Guo et al. (2023); Wang et al. (2024e); Yue et al. (2024a); Liang et al. (2024d) proposing new benchmark datasets of scientific problems but evaluating the performance of general-purpose LLMs only, and we do not include these works in our survey. Furthermore, some LLMs may belong to more than one field or modality category given our classification criteria in the paper. For instance, BioMedGPT Luo et al. (2023c) is pre-trained on biology and chemistry data jointly; GIT-Mol Liu et al. (2024a) considers the language, graph, and vision modalities simultaneously. For the sake of brevity, we introduce each of them in only one subsection.
@@ Acknowledgments
Research was supported in part by US DARPA KAIROS Program No. FA8750-19-2-1004 and INCAS Program No. HR001121C0165, National Science Foundation IIS-19-56151, and the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
@@ Appendix A Summary Tables of Scientific LLMs
GSM8K-GPT Cobbe et al. (2021b) L 6B, 175B GPT-3 GSM8K supervised fine-tuning QA, MWP
@@
MolGen Fang et al. (2024) Molecule 406M BART ZINC, NPASS sequence to sequence, molecule generation
@@ 8.9B PMC, MIMIC-III, sentence similarity estimation
MI-Zero Lu et al. (2023a) L+V BERT + histopathology figure- text-image matching classification
@@
PLIP Huang et al. (2023b) L+V GPT-2 + Twitter text-image pairs, text-image matching classification,
@@
MaCo Huang et al. (2023a) L+V BERT + MIMIC-CXR masked image modeling, classification, segmentation,

### v3

@@ 6.3 Language + Vision
Biomedical text-image pairs typically come from two sources: (1) medical reports, such as chest X-rays (e.g., MIMIC-CXR Johnson et al. (2019)) and pathology reports Huang et al. (2023); as well as (2) figure-caption pairs extracted from biomedical papers (e.g., ROCO Pelka et al. (2018) and MedICaT Subramanian et al. (2020)).
Most biomedical vision-language models exploit the CLIP architecture Radford et al. (2021), where a text encoder and an image encoder are jointly trained to map the paired text and image closer via contrastive learning (Type 3.d). The choice of the text encoder evolves from BERT Zhang et al. (2022) and GPT-2 Huang et al. (2023) to LLaMA Wu et al. (2023) and LLaMA-2 Liu et al. (2023b), while the image encoder evolves from ResNet Huang et al. (2021) to ViT Zhang et al. (2023c) and Swin Transformer Thawkar et al. (2024). MLM, masked image modeling, and text-text/image-image contrastive learning (i.e., by creating augmented views within the language/vision modality) are sometimes adopted as auxiliary pre-training tasks. Besides CLIP, other general-domain vision-language architectures, such as LLaVA Li et al. (2023a), PaLM-E Tu et al. (2024), and Gemini Saab et al. (2024), have been explored. For instance, LLaVA-Med (Type 2.d) encodes images onto several visual tokens and prepends them to text tokens as the LLM input. Evaluation tasks of these models encompass image classification, segmentation, object detection, vision QA, text-to-image/image-to-text retrieval, and report generation, the benchmarks of which include CheXpert Irvin et al. (2019), PadChest Bustos et al. (2020), SLAKE Liu et al. (2021a), etc.
@@ 6.5 Applications in Scientific Discovery
Similarly to chemistry, LLMs can automate experiments in biological and medical research. For example, CRISPR-GPT Huang et al. (2024a) augments an LLM agent with domain knowledge to enhance the design process of CRISPR-based gene-editing experiments; TrialMind Wang et al. (2024h) utilizes LLMs to extract results and synthesize clinical evidence from the literature for medical discovery. Moreover, LLMs can encode biological sequences to capture structural properties and guide protein design. For instance, ESM-1b Rives et al. (2021) and ESM-2 Lin et al. (2023b) enable accurate structure prediction of proteins without expensive and time-consuming experiments; Ferruz and Höcker (2022) fine-tune LLMs on protein families, which can generate highly divergent but still potentially functional novel sequences; He et al. (2024a) leverage an LLM for the de novo generation of SARS-CoV-2 antibodies with desired antigen-binding specificity; Hie et al. (2021) develop LLMs to evaluate the evolutionary fitness of viral variants using sequence data alone.
@@ 8 Challenges and Future Directions
This survey primarily covers LLMs in mathematics and natural sciences. We are aware that LLMs can also significantly impact social sciences by achieving remarkable performance in representative tasks Ziems et al. (2024) and serving as agents for social simulation experiments Horton (2023), but we leave the survey of these efforts as future work due to space limitations. In addition, this paper focuses on LLMs pre-trained on scientific data or augmented with domain-specific knowledge to benefit scientific discovery. There are studies Guo et al. (2023); Wang et al. (2024f); Yue et al. (2024a); Liang et al. (2024d) proposing new benchmark datasets of scientific problems but evaluating the performance of general-purpose LLMs only, and we do not include these works in our survey. Furthermore, some LLMs may belong to more than one field or modality category given our classification criteria in the paper. For instance, BioMedGPT Luo et al. (2023c) is pre-trained on biology and chemistry data jointly; GIT-Mol Liu et al. (2024a) considers the language, graph, and vision modalities simultaneously. For the sake of brevity, we introduce each of them in only one subsection.
@@ Acknowledgments
Research was supported in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, National Science Foundation IIS-19-56151, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
@@ Appendix A Summary Tables of Scientific LLMs
GSM8K-GPT Cobbe et al. (2021) L 6B, 175B GPT-3 GSM8K supervised fine-tuning QA, MWP
@@
TheoremLlama Wang et al. (2024e) L 8B LLaMA-3 Open Bootstrapped instruction tuning mathematical proof generation
Theorems
@@
MolGen Fang et al. (2024b) Molecule 406M BART ZINC, NPASS sequence to sequence, molecule generation
@@ 8.9B PMC, MIMIC-III, sentence similarity estimation
MI-Zero Lu et al. (2023) L+V BERT + histopathology figure- text-image matching classification
@@
(Biology and Medicine, Table Continued)
PLIP Huang et al. (2023) L+V GPT-2 + Twitter text-image pairs, text-image matching classification,
@@
MaCo Huang et al. (2024b) L+V BERT + MIMIC-CXR masked image modeling, classification, segmentation,
@@
(Geography, Geology, and Environmental Science, Table Continued)